Revenue 700 000
Purchases 187 000
Transport 25 000
Wages and salaries 156 000
25 percent of 525 000 interest on loan for 2 years ago
5000 rent per month

Calculate GDP using income or factor payments approach
1
Expert's answer
2020-08-24T14:53:11-0400

Solution:

GDP under factor payments approach =

"GDP = W+I+R+PR"

"W=Wages"

"I=Interests"

"R=Rents"

"PR=Business \;profits"


Therefore:

"W=156,000"

"I=25\%\times 525,000=131,250 \;for \;2years. \;1\;year \;will\;be\;\frac{131,250}{2} =65,625"

"R=5,000\;per \;month\times 12=60,000"

"PR=700,000-187,000-25,000=488,000"


Therefore GDP using the Factor payments approach will be:

"GDP=156,000+65,625+60,000+488,000"


"GDP=770,625"